Direct Materials Purchases Budget Anticipated sales for Sure Grip Tire Company were 55,000 passenger car tires and 17,000 truck tires. Rubber and steel belts are used in producing passenger car and truck tires according to the following table: Passenger Car Truck Rubber 32 lbs. per unit 75 lbs. per unit Steel belts 6 lbs. per unit 15 bs. per unit The purchase prices of rubber and steel are $3.4 and $4.5 per pound, respectively. The desired ending inventories of rubber and steel belts are 52,000 and 11,000 pounds, respectively. The estimated beginning inventories for rubber and steel belts are 61,000 and 9,000 pounds, respectively. Prepare a direct materials purchases budget for Sure Grip Tire Company for the year ended December 31, 2016. When required, enter unit prices to the nearest cent. Sure Grip Tire Compamy Direct Materials Purchases Budget For the Year Ending December 31, 2016 Rubber Steel Belts Total

Sure Grip Tire Company Direct Materials Purchases Budget For the Year Ending December 31, 2016 Rubber Steel Belts Total Pounds required for production: bs Passenger tires Truck tires bs. bs Total bs. bs. Total units purchased Unit price Total direct materials to be purchased
Answer #1

Direct material purchase budget

Rubber steel belts Total
Pounds required for production
Passenger tires 55000*32 = 1760000 55000*6 = 330000
Truck tires 17000*75 = 1275000 17000*15 = 255000
Desired ending inventory 52000 11000
Total 3087000 596000
Beginning inventory -61000 -9000
Total units purchased 3026000 587000
Unit price 3.4 4.5
Total direct materials to be purchased 10288400 2641500 12929900